OpenStudy (anonymous):

the wight of a body varies inversly as the squareof its distance from the center of the earth. if the radius of the earth is approximatly 6378km, how much would a 100 kg man weigh 2000km above the surface of the earth?

OpenStudy (anonymous):

So let the weight be a function of distance. W=k/x^2. first when x=radius of earth=6378 km W=100kg so find the constant of proportionality k. After finding k put x=2000 and find W.

OpenStudy (anonymous):

i dont understand

OpenStudy (pokemon23):

hero btw i got the glasses?

hero (hero):

Thank goodness

OpenStudy (pokemon23):

I thought you were going to say that

OpenStudy (anonymous):

Ok look W is proportional to 1/x^2. So then w=k/x^2.I just put a constant of proportionality k.Understood?

OpenStudy (anonymous):

anyone understands what he meant to explain me in a simple way

OpenStudy (anonymous):

\[w=k/d ^{2}\] given radius=6378; w=100kg\[100=k/6378^{2}\]\[100=k/40678884\]\[k=100(40678884)\]\[k=4067888400\]then proceed on finding W. since we have the value of k d= 6378(from the center to the surface) + 2000(above the surface) \[w=4067888400/8378\]\[w=4855440917 kg.\]

OpenStudy (anonymous):

is that what you ment before shankvee\

OpenStudy (anonymous):

oh i forgot square the denominator first before dividing it. .

OpenStudy (anonymous):

that's 57.95 kg

OpenStudy (anonymous):

what did you do to get 57.95

OpenStudy (anonymous):

square 6378, then divide to the numerator. .

OpenStudy (anonymous):

you lost me now can you start from where you forgot to do that

OpenStudy (anonymous):

please can you help me Jerwyn gayo